OpenStudy (anonymous):

can someone simplify the squareroot of 50x^7y^4?

OpenStudy (anonymous):

=5*x^3*y^2*sqrt(2x)

OpenStudy (anonymous):

\[5x ^{3}y ^{2}\sqrt{2y}\]

OpenStudy (anonymous):

\[\sqrt{50c^7y^4}=\sqrt{25\times 2\times x^6\times y^4}=\sqrt{25}\sqrt{x^6}\sqrt{y^4}\sqrt{2x}\] \[=5x^3y^2\sqrt{2x}\]